OpenStudy (anonymous):

PLEASE HELP ME !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!A cubical box is 729 cubic centimeters. What is the length of one side of the box?the answers are A. 27 cm B. 18 cm C. 9 cm D. 3 cm

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

Hola, ¿cómo estás hoy? ¿Puedo ayudarle? :) English or espanol? @donavan

OpenStudy (anonymous):

Si, si!

OpenStudy (anonymous):

english

OpenStudy (anonymous):

oh ok hi ok wat is te problam???

OpenStudy (anonymous):

i can figer out how to do this

OpenStudy (anonymous):

pero, eso very easy ok i show u how ok?

OpenStudy (anonymous):

i show you how ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

please

OpenStudy (anonymous):

ok that is the whole problam????

OpenStudy (anonymous):

yes that is the whole problem

OpenStudy (anonymous):

do u kno how to divide you divide ok you do tat to get it ok and wen u divide the 729 divide them to all the choices do you get yet????

OpenStudy (jdoe0001):

so whatever happened to \(\huge \bf \sqrt[3]{729} \ \ ?\)

OpenStudy (anonymous):

what do i divide

OpenStudy (jdoe0001):

is not division, is a radical

OpenStudy (anonymous):

ok ok @jdoe0001 ya he kno how to do it ha you can win dis one easy with him you got it ok follow jdoe ;) glad to help u papi

OpenStudy (jdoe0001):

\(\bf \huge \sqrt[2]{4} \implies \sqrt{4} \implies \sqrt{2^2} \implies 2\)

OpenStudy (anonymous):

so how would i do this

OpenStudy (jdoe0001):

got a calculator?

OpenStudy (jdoe0001):

or lemme try google's calc :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont know what that is

OpenStudy (jdoe0001):

well, that's google's calculator :)

OpenStudy (anonymous):

ok on there

OpenStudy (anonymous):

what know

OpenStudy (anonymous):

ok i got 2

OpenStudy (jdoe0001):

\(\bf \huge \sqrt[3]{729} \implies \sqrt[3]{9\times 9 \times 9}\implies \sqrt[\color{red}{3}]{9^\color{red}{3}}\implies 9 \)

OpenStudy (jdoe0001):

is a cubical box, all its sides are the same value, so length = 9 width = 9 height = 9 9x9x9 = 729\(cm^3\)

OpenStudy (jdoe0001):